OpenStudy (anonymous):

The line through the origin perpendicular to the line x-3y=9 (write an equation of the line described)

OpenStudy (inkyvoyd):

you have f(0)=0 and you want to find the line perpendicular to x-3y=9 Note that perpendicular lines have slopes that are -1/m, if m is the slope of the original line.

OpenStudy (anonymous):

thanks, I am having trouble working this out, can you help me?

OpenStudy (inkyvoyd):

write the equation for the line you're given in terms of y, so it looks like y=blah blah

OpenStudy (inkyvoyd):

(move everything by y to the right side of the equation)

OpenStudy (anonymous):

this is what I got -3y=9-x

OpenStudy (inkyvoyd):

divide both sides by -3

OpenStudy (inkyvoyd):

y=x/3-3

OpenStudy (inkyvoyd):

what is the slope?

OpenStudy (anonymous):

i got 1/3

OpenStudy (inkyvoyd):

ok, so if m=1/3 what is -1/m?

OpenStudy (anonymous):

im not sure... that's what im confused on

OpenStudy (inkyvoyd):